OpenStudy (anonymous):

Find the derivative: y = cosh^-1(x)^1/2

OpenStudy (abb0t):

[cosh\(^{-1}\)x]\(^\frac{1}{2}\) ????

OpenStudy (mandre):

\[\frac{ d }{ dx } \cosh ^{-1}(x ^{\frac{ 1 }{ 2 }})= \frac{ 1 }{ \sqrt{x-1} }.\frac{ 1 }{ 2 }x ^{-\frac{ 1 }{ 2 }}=\frac{ 1 }{2\sqrt{x} \sqrt{x-1} }=\frac{ 1 }{2 \sqrt{x ^{2}-x} }\]

OpenStudy (anonymous):

To @abb0t yes, [cosh−1x]^(1/2)